OpenStudy (anonymous):

Solve. 10x^4-11x^2+3=0

OpenStudy (akshay_budhkar):

10x^4-5x^2-6x^2+3=0 5x^2(2x^2-1)-3(2x^2-1)=0 (5x^2-3)(2x^2-1)=0

OpenStudy (akshay_budhkar):

therefore x^2=3/5 or 1/2 so x is =3/5,-3/5,1/2,-1/2

OpenStudy (anonymous):

That made things more understandable. I almost forgot how to do this. Thank you!

OpenStudy (akshay_budhkar):

np yw :D